OpenStudy (anonymous):

please help 2a-6a =10

OpenStudy (anonymous):

-5/2

OpenStudy (agreene):

2a-6a=10 (2-6)a=10 -4a=10 divide by -4 a=10/-4 simplify by taking 2 out of top and bottom. a=-5/2